OpenStudy (anonymous):

a culture started with 2000 bacteria. after 2 hour , it grew to 2,400 bacteria. predict how many bacteria will be present after 10 hours. round your answer to the nearest whole number. p=Ae^kt

OpenStudy (anonymous):

For t = 0, A = 2000. So, the population function is: \( p = 2000 e^{kt} \) Plug in t = 2 to get this equation: \(2400 = 2000e^{2k} \) Solve for k. After that, plug in t = 10 and solve for p :-)

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

You will have your answer.

OpenStudy (anonymous):

not yet! 1 moes step?

OpenStudy (anonymous):

Rounding? That should be easy, no?

OpenStudy (anonymous):

nope! i you help?

OpenStudy (anonymous):

Huh? And, after you solve for k, k is just a constant. You will have something like: \[ P(t) = 2000e^{kt} \]Where you will know k. So, solve the equation: \[ P(t) = 2000e^{10k} \]with the k being known.

OpenStudy (anonymous):

k=09115 is correct?

OpenStudy (anonymous):

I will solve it, give me a min.

OpenStudy (anonymous):

Got k = 0.0911607783969

OpenStudy (anonymous):

nope i think wrong how come ?

OpenStudy (anonymous):

2400 = 2000e^(2k) 24/20 = e^(2k) ln(24/20) = 2k k = (ln(24/20))/2 :-)

OpenStudy (anonymous):

here : solve for k : 2.400=2.000 e^k*2 1.2=e ^k*2 In(1.2)=In(e^k*2) 0.1823=k*2 and after this one at 10 hr t=10 i don't got this step!?

OpenStudy (anonymous):

Ok, you are having rounding issues with your k. Do not round it. Afterwards: P(10) = 2000e^(0.0911607783969*10)

OpenStudy (anonymous):

so what is the end answer?

OpenStudy (anonymous):

4977, somewhere near that.

OpenStudy (anonymous):

how u got 0.0911....can you help me that step because i got problem with that that y i dont know?

OpenStudy (anonymous):

You probably read wrong on your calculator. It should be 0.01823. You are on the right track.

OpenStudy (anonymous):

no.. ur wrong

OpenStudy (anonymous):

How much is 0.18 divided by 2? :-)